# Radiative decays of $P$-wave bottom baryons from light-cone sum rules

### v1

@@ III.5 The $[\mathbf{\bar{3}}_{F},2,1,\rho]$ doublet
Doublets Baryon       Mass   Difference    Decay channels   Decay width     T.R.W.    T.S.W. Tan:2023opd
($j^{P}$) (GeV) (MeV) (keV) (keV) (MeV)
  $[\mathbf{\bar{3}}_{F},0,1,\rho]$   $\Xi_{b}({1\over 2}^{-})$ $6.10^{+0.08}_{-0.08}$ $\Xi_{b}^{0}({1\over 2}^{-})\to\Xi_{b}^{0}\gamma$ $440^{+390}_{-440}$ $1200^{+1000}_{-1200}$ $12000^{+10000}_{-~{}9000}$
$\Xi_{b}^{0}({1\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $720^{+640}_{-720}$
$\Xi_{b}^{-}({1\over 2}^{-})\to\Xi_{b}^{-}\gamma$ $4.4^{+1.5}_{-4.0}$ $12.9^{+~{}8.4}_{-12.4}$
$\Xi_{b}^{-}({1\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $8.5^{+6.9}_{-8.4}$
  $[\mathbf{\bar{3}}_{F},1,1,\rho]$   $\Xi_{b}({1\over 2}^{-})$ $6.09^{+0.13}_{-0.12}$ $7\pm 3$ $\Xi_{b}^{0}({1\over 2}^{-})\to\Xi_{b}^{0}\gamma$ $41^{+78}_{-41}$ $46^{+100}_{-~{}46}$ $3.7^{+9.5}_{-3.7}$
$\Xi_{b}^{0}({1\over 2}^{-})\to\Xi_{b}^{\prime 0}\gamma$ $3.5^{+17.9}_{-~{}3.5}$
$\Xi_{b}^{0}({1\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $1.2^{+8.1}_{-1.2}$
$\Xi_{b}^{-}({1\over 2}^{-})\to\Xi_{b}^{-}\gamma$ $97^{+200}_{-~{}91}$ $97^{+200}_{-~{}91}$
$\Xi_{b}^{-}({1\over 2}^{-})\to\Xi_{b}^{\prime-}\gamma$ $0.01^{+0.03}_{-0.01}$
$\Xi_{b}^{-}({1\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $0.01^{+0.08}_{-0.01}$
  $\Xi_{b}({3\over 2}^{-})$ $6.10^{+0.13}_{-0.12}$ $\Xi_{b}^{0}({3\over 2}^{-})\to\Xi_{b}^{0}\gamma$ $85^{+160}_{-~{}85}$ $87^{+170}_{-~{}87}$ $0.64^{+1.70}_{-0.64}$
$\Xi_{b}^{0}({3\over 2}^{-})\to\Xi_{b}^{\prime 0}\gamma$ $1.7^{+7.9}_{-1.7}$
$\Xi_{b}^{0}({3\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $0.84^{+4.89}_{-0.84}$
$\Xi_{b}^{-}({3\over 2}^{-})\to\Xi_{b}^{-}\gamma$ $220^{+420}_{-210}$ $220^{+430}_{-210}$
$\Xi_{b}^{-}({3\over 2}^{-})\to\Xi_{b}^{\prime-}\gamma$ $1.5^{+6.8}_{-1.5}$
$\Xi_{b}^{-}({3\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $0.75^{+5.19}_{-0.75}$
  $[\mathbf{\bar{3}}_{F},2,1,\rho]$   $\Xi_{b}({3\over 2}^{-})$ $6.10^{+0.15}_{-0.10}$ $14\pm 7$ $\Xi_{b}^{0}({3\over 2}^{-})\to\Xi_{b}^{\prime 0}\gamma$ $140^{+510}_{-140}$ $150^{+550}_{-150}$ $1.4^{+11.6}_{-~{}1.4}$
$\Xi_{b}^{0}({3\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $7.8^{+35.5}_{-~{}7.8}$
$\Xi_{b}^{-}({3\over 2}^{-})\to\Xi_{b}^{\prime-}\gamma$ $0.98^{+3.52}_{-0.98}$ $1.0^{+3.8}_{-1.0}$
$\Xi_{b}^{-}({3\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $0.10^{+0.31}_{-0.10}$
  $\Xi_{b}({5\over 2}^{-})$ $6.11^{+0.15}_{-0.10}$ $\Xi_{b}^{0}({5\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $19^{+74}_{-19}$ $19^{+74}_{-19}$ $1.0^{+7.4}_{-1.0}$
$\Xi_{b}^{-}({5\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $8.4^{+30.9}_{-~{}8.3}$ $8.4^{+30.9}_{-~{}8.3}$
  $[\mathbf{\bar{3}}_{F},1,0,\lambda]$   $\Xi_{b}({1\over 2}^{-})$ $6.10^{+0.20}_{-0.10}$ $4\pm 2$ $\Xi_{b}^{0}({1\over 2}^{-})\to\Xi_{b}^{\prime 0}\gamma$ $1200^{+11000}_{-~{}1200}$ $1800^{+17000}_{-~{}1800}$ $2800^{+8600}_{-2800}$
$\Xi_{b}^{0}({1\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $550^{+6400}_{-~{}550}$
$\Xi_{b}^{-}({1\over 2}^{-})\to\Xi_{b}^{\prime-}\gamma$ $3.8^{+36.2}_{-~{}3.8}$ $5.5^{+57.0}_{-~{}5.5}$
$\Xi_{b}^{-}({1\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $1.7^{+20.9}_{-~{}1.7}$
  $\Xi_{b}({3\over 2}^{-})$ $6.10^{+0.20}_{-0.10}$ $\Xi_{b}^{0}({3\over 2}^{-})\to\Xi_{b}^{\prime 0}\gamma$ $660^{+5800}_{-~{}660}$ $700^{+6200}_{-~{}700}$ $330^{+2020}_{-~{}330}$
$\Xi_{b}^{0}({3\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $38^{+430}_{-~{}38}$
$\Xi_{b}^{-}({3\over 2}^{-})\to\Xi_{b}^{\prime-}\gamma$ $2.1^{+20.6}_{-~{}2.1}$ $2.2^{+21.9}_{-~{}2.2}$
$\Xi_{b}^{-}({3\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $0.12^{+1.34}_{-0.12}$

### v3

@@ III Decays of flavor $\mathbf{\bar{3}}_{F}$ baryons
Doublets Baryon       Mass   Difference    Decay channels  Coupling constants   Decay width     T.R.W.
($j^{P}$) (GeV) (MeV) ($\rm GeV^{-1}$) (keV) (keV)
  $[\mathbf{\bar{3}}_{F},0,1,\rho]$   $\Xi_{b}({1\over 2}^{-})$ $6.10^{+0.08}_{-0.08}$ $\Xi_{b}^{0}({1\over 2}^{-})\to\Xi_{b}^{0}\gamma$ $1.00^{+0.38}_{-0.30}$ $440^{+390}_{-440}$ $1200^{+1000}_{-1200}$
$\Xi_{b}^{0}({1\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $1.10^{+0.41}_{-0.31}$ $720^{+640}_{-720}$
$\Xi_{b}^{-}({1\over 2}^{-})\to\Xi_{b}^{-}\gamma$ $0.10^{+0.01}_{-0.02}$ $4.4^{+1.5}_{-4.0}$ $12.9^{+~{}8.4}_{-12.4}$
$\Xi_{b}^{-}({1\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $0.12^{+0.01}_{-0.03}$ $8.5^{+6.9}_{-8.4}$
  $[\mathbf{\bar{3}}_{F},1,1,\rho]$   $\Xi_{b}({1\over 2}^{-})$ $6.09^{+0.13}_{-0.12}$ $7\pm 3$ $\Xi_{b}^{0}({1\over 2}^{-})\to\Xi_{b}^{0}\gamma$ $0.13^{+0.04}_{-0.06}$ $41^{+78}_{-41}$ $46^{+100}_{-~{}46}$
$\Xi_{b}^{0}({1\over 2}^{-})\to\Xi_{b}^{\prime 0}\gamma$ $0.10^{+0.03}_{-0.05}$ $3.5^{+17.9}_{-~{}3.5}$
$\Xi_{b}^{0}({1\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $0.052^{+0.020}_{-0.021}$ $1.2^{+8.1}_{-1.2}$
$\Xi_{b}^{-}({1\over 2}^{-})\to\Xi_{b}^{-}\gamma$ $0.20^{+0.07}_{-0.05}$ $97^{+200}_{-~{}91}$ $97^{+200}_{-~{}91}$
$\Xi_{b}^{-}({1\over 2}^{-})\to\Xi_{b}^{\prime-}\gamma$ $0.004^{+0.003}_{-0.004}$ $0.01^{+0.03}_{-0.01}$
$\Xi_{b}^{-}({1\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $0.005^{+0.002}_{-0.002}$ $0.01^{+0.08}_{-0.01}$
  $\Xi_{b}({3\over 2}^{-})$ $6.10^{+0.13}_{-0.12}$ $\Xi_{b}^{0}({3\over 2}^{-})\to\Xi_{b}^{0}\gamma$ $0.18^{+0.06}_{-0.07}$ $85^{+160}_{-~{}85}$ $87^{+170}_{-~{}87}$
$\Xi_{b}^{0}({3\over 2}^{-})\to\Xi_{b}^{\prime 0}\gamma$ $0.064^{+0.026}_{-0.026}$ $1.7^{+7.9}_{-1.7}$
$\Xi_{b}^{0}({3\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $0.074^{+0.030}_{-0.031}$ $0.84^{+4.89}_{-0.84}$
$\Xi_{b}^{-}({3\over 2}^{-})\to\Xi_{b}^{-}\gamma$ $0.29^{+0.10}_{-0.10}$ $220^{+420}_{-210}$ $220^{+430}_{-210}$
$\Xi_{b}^{-}({3\over 2}^{-})\to\Xi_{b}^{\prime-}\gamma$ $0.006^{+0.002}_{-0.002}$ $1.5^{+6.8}_{-1.5}$
$\Xi_{b}^{-}({3\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $0.007^{+0.002}_{-0.002}$ $0.75^{+5.19}_{-0.75}$
  $[\mathbf{\bar{3}}_{F},2,1,\rho]$   $\Xi_{b}({3\over 2}^{-})$ $6.10^{+0.15}_{-0.10}$ $14\pm 7$ $\Xi_{b}^{0}({3\over 2}^{-})\to\Xi_{b}^{\prime 0}\gamma$ $0.57^{+0.26}_{-0.20}$ $140^{+510}_{-140}$ $150^{+550}_{-150}$
$\Xi_{b}^{0}({3\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $0.22^{+0.13}_{-0.10}$ $7.8^{+35.5}_{-~{}7.8}$
$\Xi_{b}^{-}({3\over 2}^{-})\to\Xi_{b}^{\prime-}\gamma$ $0.048^{+0.014}_{-0.012}$ $0.98^{+3.52}_{-0.98}$ $1.0^{+3.8}_{-1.0}$
$\Xi_{b}^{-}({3\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $0.021^{+0.010}_{-0.012}$ $0.10^{+0.31}_{-0.10}$
  $\Xi_{b}({5\over 2}^{-})$ $6.11^{+0.15}_{-0.10}$ $\Xi_{b}^{0}({5\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $0.18^{+0.10}_{-0.10}$ $19^{+74}_{-19}$ $19^{+74}_{-19}$
$\Xi_{b}^{-}({5\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $0.014^{+0.003}_{-0.003}$ $8.4^{+30.9}_{-~{}8.3}$ $8.4^{+30.9}_{-~{}8.3}$
  $[\mathbf{\bar{3}}_{F},1,0,\lambda]$   $\Xi_{b}({1\over 2}^{-})$ $6.10^{+0.20}_{-0.10}$ $4\pm 2$ $\Xi_{b}^{0}({1\over 2}^{-})\to\Xi_{b}^{\prime 0}\gamma$ $1.70^{+0.58}_{-0.47}$ $1200^{+11000}_{-~{}1200}$ $1800^{+17000}_{-~{}1800}$
$\Xi_{b}^{0}({1\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $0.97^{+0.27}_{-0.24}$ $550^{+6400}_{-~{}550}$
$\Xi_{b}^{-}({1\over 2}^{-})\to\Xi_{b}^{\prime-}\gamma$ $0.095^{+0.061}_{-0.070}$ $3.8^{+36.2}_{-~{}3.8}$ $5.5^{+57.0}_{-~{}5.5}$
$\Xi_{b}^{-}({1\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $0.055^{+0.037}_{-0.040}$ $1.7^{+20.9}_{-~{}1.7}$
  $\Xi_{b}({3\over 2}^{-})$ $6.10^{+0.20}_{-0.10}$ $\Xi_{b}^{0}({3\over 2}^{-})\to\Xi_{b}^{\prime 0}\gamma$ $1.20^{+0.41}_{-0.30}$ $660^{+5800}_{-~{}660}$ $700^{+6200}_{-~{}700}$
$\Xi_{b}^{0}({3\over 2}^{-})\to\Xi_{b}^{*0}\gamma$ $0.47^{+0.17}_{-0.11}$ $38^{+430}_{-~{}38}$
$\Xi_{b}^{-}({3\over 2}^{-})\to\Xi_{b}^{\prime-}\gamma$ $0.067^{+0.049}_{-0.049}$ $2.1^{+20.6}_{-~{}2.1}$ $2.2^{+21.9}_{-~{}2.2}$
$\Xi_{b}^{-}({3\over 2}^{-})\to\Xi_{b}^{*-}\gamma$ $0.026^{+0.017}_{-0.018}$ $0.12^{+1.34}_{-0.12}$